An experiment consists of rolling 2 dice and observing the number that lands on the uppermost surface. Find the probability that the one of the numbers is a 5, given that the sum of the numbers is even.

Explain and answer in decimals please, thank you.
Answer #1

P(one of the number is 5 | sum of the numbers is even)

= P(one of the number is 5 and sum of the numbers is even)/P(sum of the numbers is even)

= 4/18 = 0.2222